# Automated Contrastive Learning Strategy Search for Time Series

### v1

@@ Abstract.
In recent years, Contrastive Learning (CL) has become a predominant representation learning paradigm for time series. Most existing methods in the literature focus on manually building specific Contrastive Learning Strategies (CLS) by human heuristics for certain datasets and tasks. However, manually developing CLS usually require excessive prior knowledge about the datasets and tasks, e.g., professional cognition of the medical time series in healthcare, as well as huge human labor and massive experiments to determine the detailed learning configurations. In this paper, we present an Automated Machine Learning (AutoML) practice at Microsoft, which automatically learns to contrastively learn representations for various time series datasets and tasks, namely Automated Contrastive Learning (AutoCL). We first construct a principled universal search space of size over $3\times 10^{12}$, covering data augmentation, embedding transformation, contrastive pair construction and contrastive losses. Further, we introduce an efficient reinforcement learning algorithm, which optimizes CLS from the performance on the validation tasks, to obtain more effective CLS within the space. Experimental results on various real-world tasks and datasets demonstrate that AutoCL could automatically find the suitable CLS for a given dataset and task. From the candidate CLS found by AutoCL on several public datasets/tasks, we compose a transferable Generally Good Strategy (GGS), which has a strong performance for other datasets. We also provide empirical analysis as a guidance for future design of CLS.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Time series is a sequence of data points recorded chronologically, which has been collected and analyzed for various real-world applications such as clinical diagnosis (Eldele et al., 2021; Zhang et al., 2022b), electricity usage forecasting (Zhou et al., 2021; Yue et al., 2022) and web traffic monitoring (Ren et al., 2019). A paramount challenge in the field of time series representation learning lies in developing an efficacious encoder capable of deriving informative embeddings. These embeddings should be readily applicable to a variety of downstream tasks, such as classification and forecasting, thereby enhancing the model’s utility and performance in predictive analytics.
@@
A predominant representation learning paradigm is Contrastive Learning (CL) (Liu et al., 2021), of which the core idea is to train models by pulling the positive embedding pairs closer and pushing the negative embedding pairs far apart. CL has been widely explored in domains such as computer vision (Ozbulak et al., 2023), natural language processing (Zhang et al., 2022a) and graph mining (Wu et al., 2021). In recent years, CL has been introduced into the time series domain (Zhang et al., 2023). The preponderance of existing research in CL as applied to time series data primarily emphasizes the development of sophisticated data augmentation techniques and the construction of contrastive pairs. For instance, TS-TCC (Eldele et al., 2021) employs jittering and scaling techniques to produce varied perspectives of the input time series. These generated perspectives, originating from the same data sample, are then treated as positive instances within the model’s training regime. TS2Vec (Yue et al., 2022) uses random cropping to generate positive pairs and treat other data points in the context as negative samples. TF-C (Zhang et al., 2022b) augments the input via time-domain and frequency-domain masking, and embeddings from different samples are treated as negative pairs.
@@
Despite the effectiveness of these approaches, the majority are tailored with manual construction of Contrastive Learning Strategies (CLS) specific to particular datasets and tasks, representing distinct examples within the broader spectrum of general CLS space, which could contain millions even trillions of choices. Crafting CLS for specific tasks or datasets in such a huge space through heuristic methods could be challenging, often demanding extensive domain expertise and considerable manual effort through numerous trial-and-error iterations. A natural question arises that is it possible to build a system to automatically find a suitable CLS for a given task? A pertinent question emerges: can we identify a CLS that is generally effective across a dynamic range of tasks and datasets?
@@
Aiming at automating the process of machine learning, Automated Machine Learning (AutoML) (He et al., 2021) has become popular in recent years. In the field of CL, some works optimize data augmentation strategy on image (Tian et al., 2020) and graph (You et al., 2021) data. For time series, a recent work InfoTS (Luo et al., 2023) also searches the data augmentation based on information theory which yet does not consider the final goal of representation learning for downstream tasks. Moreover, these studies predominantly focus on data augmentation, they often overlook other crucial dimensions of CL, such as loss functions, the construction of contrastive pairs, and embedding transformations, which are typically crafted by human intervention. Specifically, a CLS encompasses diverse loss functions (e.g., InfoNCE (Oord et al., 2018) and Triplet Loss (Hoffer and Ailon, 2015)) and metric functions (e.g., dot product, cosine similarity, negative Euclidean distance) that affect representation learning quality (Liu et al., 2021). Contrastive pair construction also plays a key role to distinguish sample-level characteristics or temporal dependencies (Oord et al., 2018). Embedding techniques, including augmentation (Wang et al., 2019) and normalization (Wang and Isola, 2020), further influences the pretrained model performance. Few works have been conducted on jointly considering all these aspects of the contrastive learning on time series.
@@
To reduce the burden of excessive prior domain knowledge and massive manual trials of CL on time series for specific tasks, in this paper, we present an Automated Contrastive Learning (AutoCL) framework at Microsoft, which aims at learning to contrastively learn the representations for various time series tasks. Specifically, we first construct a comprehensive solution space of CLS, which not only covers the most important configurable dimensions of CL, but also considers the modest range of options for each dimension. The whole space is comprised of dimensions including data augmentations, embedding transformations, contrastive pair construction and contrastive learning objectives, with up to $10^{12}$ level of options in total. To efficiently target the suitable CLS in such huge solution space, we further introduce a reinforcement learning algorithm which directly optimizes the CL performance on downstream tasks in validation datasets. Compared with the recent automation-based CL method InfoTS (Luo et al., 2023), our method takes the two-fold advantage from more comprehensive solution space and the direct optimization approach on CL, while InfoTS only searches data augmentation configurations based on task-agnostic criteria, which does not consider the intrinsic connection between CLS and the downstream task, thus may result in targeting an improper CLS.
@@
We empirically evaluate our AutoCL framework and the derived CLS over three practical downstream tasks including classification, forecasting and anomaly detection on real-world time series datasets. We first directly apply AutoCL over each task to obtain its suitable CLS, and the results of which could significantly outperform existing CL works, demonstrating the superiority of AutoCL. Then we derive a Generally Good Strategy (GGS) based on the candidate CLS found by AutoCL from several tasks, and test GGS on all tasks. The results show that GGS could achieve remarkable performance for all tasks, showing its strong transferability across tasks and datasets. We also provide extensive empirical analysis on the relation between the candidate CLS and the model’s performance on downstream tasks, hoping to provide guidance for the future design of CLS. We also tested our derived solution in a real-world deployed application of epilepsy seizure detection task in Shanghai Ruijin hospital in China, which has demonstrated the general good performance on the realistic platform.
The major contributions of the paper are summarized as follows:
@@
  • We introduce AutoCL to automatically derive suitable CLS for various tasks and datasets on time series. A comprehensive solution space covering the most critical dimensions of CL with an efficient reinforcement learning algorithm have been proposed.
@@ 2. Methodology
In this section, we first formulate the problem of the automated contrastive learning strategy search, and then introduce a method called AutoCL for the problem. We will elaborate the two components of AutoCL: the solution space and the search algorithm.
@@ 2.1. Problem Formulation
Denote $\mathcal{D}$ as a time series dataset comprised of training $\mathcal{D}_{\text{train}}$, validation $\mathcal{D}_{\text{val}}$ and test $\mathcal{D}_{\text{test}}$ sets; $\mathcal{T}$ as a downstream task along with evaluation metrics $\mathcal{M}$; and $f^{E}$ as the encoder which extracts embeddings of the input time series. The problem of automated Contrastive Learning Strategy (CLS) search has two sub-problem: (1) define a solution space $\mathcal{A}$; (2) build a search algorithm $f$ to find a suitable CLS $A\in\mathcal{A}$ based on the validation performance on $\mathcal{D}_{\text{val}}$ regarding $\mathcal{M}$, such that after pre-training with $A$ on $\mathcal{D}_{\text{train}}$, $f^{E}$ will have a good performance on $\mathcal{D}_{\text{test}}$.
@@ 2.2. Solution Space
Based on the studies in the literature and our own experiences, we use the following principles to define the space (You et al., 2020): (1) encompasses the key dimensions of CLS, incorporating insights from prior human-designed strategies; (2) accounts for a moderate spectrum of choices within each dimension.
An illustration of a CLS is shown in Figure 1. The essence of CL is to extract semantic-preserving embeddings invariant to small data perturbations (Tian et al., 2020), and thus data augmentation is the cornerstone dimension of CLS. Another commonly used practice to enhance the robustness of representation is embedding transformations, including embedding augmentation (Wang et al., 2019) and normalization. There are various aspects to construct positive and negative embeddings pairs, where different aspects reflect different characteristics of data. Different tasks might place emphasis on different aspects. Finally, different similarity metric functions and different types of contrastive learning losses could have divergent impact on the learned representations. In summary, we consider four dimensions: data augmentations, embedding transformations, contrastive pair construction and contrastive losses. A summary of the dimensions, sub-dimensions and their options is shown in Table 1, and we describe their details in the following content.
@@
Frequency Masking (Zhang et al., 2022b)
@@
Similarity Functions Dot, Cosine, Distance
@@
Data Augmentations. Data augmentations transform the input data into different but related views, which are the cornerstones of a CLS. We consider 6 most commonly adopted data augmentations as the sub-dimensions, including resizing (length) (Mehari and Strodthoff, 2022), rescaling (amplitude) (Han et al., 2021), jittering (Eldele et al., 2021), point masking (Zerveas et al., 2021), frequency masking (Zhang et al., 2022b), and random cropping (Yue et al., 2022). Each augmentation is associated with a parameter $p\in[0,1]$. Let $\mathbf{x}\in\mathbb{R}^{T\times c}$ be an input time series, where $T$ and $c$ are the temporal length and the number of variables. For point masking, $p$ refers to the ratio of the input data points to be masked: $\mathbf{x}=\mathbf{x}\odot\mathbf{m}$, where $\mathbf{m}[t]\sim Bernoulli(1-p)$. For jittering, $p$ is the standard deviation of the Gaussian noise: $\mathbf{x}=\mathbf{x}+\mathcal{N}(0,p)$. For more details of each augmentation, please refer to the corresponding paper. We discretize the value of $p$ into 11 choices: $[0.0,0.1,\cdots,0.8,0.9,0.95]$. For the first value, we use $0.0$ to represent the setting where a specific data augmentation is not enabled. For the last value, we do not use $1.0$ since it is meaningless for some data augmentations in practice. For example, in point masking, $p=1.0$ means masking out all the input data.
Additionally, the order of applying data augmentations also influences the learned embeddings (Cubuk et al., 2019). For instance, the outcome of applying point masking followed by random cropping differs from the sequence of first random cropping and then point masking. However, the ordering of other data augmentations, such as resizing and rescaling, does not have a significant influence on the data. Therefore, we design 5 different orders of applying data augmentations. For details, please refer to Appendix. In summary, there are totally $5\times 11^{6}$ options for data augmentations.
Embedding Transformations. Embedding augmentations (Wang et al., 2019) and normalization (Wang and Isola, 2020) have been explored and proven useful in the literature. For embedding augmentations, we consider embedding jittering (Wang et al., 2019) and embedding masking as the sub-dimensions, which are similar to the jittering and point masking in data augmentation operation. These two embedding augmentations are also parameterized by a parameter $p\in[0,1]$ whose value is also discretize $p$ into 11 values. For normalization, we consider to use no norms, $l_{2}$ norm, and layer norm (Ba et al., 2016) as the options. In summary, there are $3\times 11^{2}$ options for embedding transformations.
@@
Let $\mathbf{h}_{1},\mathbf{h}_{2}\in\mathbb{R}^{T\times d}$ be the embeddings of the two views $\mathbf{x}_{1}$, $\mathbf{x}_{2}$ of an input $\mathbf{x}$, where $T$ and $d$ are the length and dimension size. The objective of the instance contrast is to distinguish embeddings of two different time series instances $\mathbf{x},\mathbf{x}^{\prime}$ by pulling $\mathbf{h}_{1}$ and $\mathbf{h}_{2}$ (as well as $\mathbf{h}_{1}^{\prime}$ and $\mathbf{h}_{2}^{\prime}$) closer and pushing $\mathbf{h}_{1},\mathbf{h}_{2}$ far away from $\mathbf{h}_{1}^{\prime},\mathbf{h}_{2}^{\prime}$. The temporal contrast aims to preserve the local temporal information. For a time step $t$, it pulls $\mathbf{h}_{1}[t]$ and $\mathbf{h}_{2}[t]$ closer and optionally also $\mathbf{h}_{1}[t]$ and its temporally adjacent embeddings $\mathbf{h}_{1}[t-1],\mathbf{h}_{1}[t+1]$ closer. It pushes $\mathbf{h}_{1}[t]$ away from $\mathbf{h}_{1}[t^{\prime}]$, where $t^{\prime}\notin[t-1,t+1]$. The above two contrasts might be sufficient for short time series (e.g., $T<100$), however, they are incapable of capturing multi-scale dynamics of long time series (e.g., $T>1000$) (Yue et al., 2022). We follow (Yue et al., 2022) and apply hierarchical pooling over $\mathbf{h}_{1},\mathbf{h}_{2}$ to obtain multi-scale embeddings $\mathbf{h}_{1}^{(s)},\mathbf{h}_{2}^{(s)}\in\mathbb{R}^{T_{s}\times d}$, where $s\in[1,\cdots,S]$ is the scale index and $T_{s}$ is the length. The hierarchical pooling is associated with two parameters: pooling operator (average or max pooling) and kernel size ($[0,2,3,5]$). The above instance contrast and temporal contrast are applied for all the scales. In addition to the instance and temporal contrasts, cross-scale contrast (Bachman et al., 2019) could further improve the cross-scale consistency. Specifically, if $\mathbf{h}^{(s+1)}[t_{s+1}]$ is pooled from $\mathbf{h}^{(s)}[t_{s}]$, then these two embeddings should be pulled closer, otherwise, they should be pushed away from each other.
@@
In practice, one must choose at least one contrastive aspect to perform CL. Based on the studies in the literature and our own experiences, the instance contrast is the most indispensable aspect, and therefore it is always included. As a result, there are 2=3-1 contrast aspects, 2 pooling operators, 4 kernel sizes and 2 options (whether to include temporal adjacent embeddings) for the temporal contrast, which means we have 32 options in total.
@@
Contrastive Losses. Different types of contrastive losses behave differently in practice. For example, InfoNCE (Oord et al., 2018) guide the model’s learning by maximizing the probabilities of the positive pairs, and Triplet loss (Hoffer and Ailon, 2015) directly maximize the similarity score of the positive pairs. A contrastive loss, e.g., InfoNCE, is usually comprised of a similarity function, measuring the similarity between two embeddings, and a temperature parameter (Oord et al., 2018), controling the magnitude of similarly scores. For the similarity functions, we consider dot product, cosine and negative Euclidean distance. For the temperature parameter, we consider $[10^{-2},10^{-1},10^{0},10^{1},10^{2}]$. In summary, there are 2 types of losses, 3 similarity functions and 5 temperatures, resulting in 30 total options.

### v2

@@ Abstract.
In recent years, Contrastive Learning (CL) has become a predominant representation learning paradigm for time series. Most existing methods manually build specific CL Strategies (CLS) by human heuristics for certain datasets and tasks. However, manually developing CLS usually requires excessive prior knowledge about the data, and massive experiments to determine the detailed CL configurations. In this paper, we present an Automated Machine Learning (AutoML) practice at Microsoft, which automatically learns CLS for time series datasets and tasks, namely Automated Contrastive Learning (AutoCL). We first construct a principled search space of size over $3\times 10^{12}$, covering data augmentation, embedding transformation, contrastive pair construction, and contrastive losses. Further, we introduce an efficient reinforcement learning algorithm, which optimizes CLS from the performance on the validation tasks, to obtain effective CLS within the space. Experimental results on various real-world datasets demonstrate that AutoCL could automatically find the suitable CLS for the given dataset and task. From the candidate CLS found by AutoCL on several public datasets/tasks, we compose a transferable Generally Good Strategy (GGS), which has a strong performance for other datasets. We also provide empirical analysis as a guide for the future design of CLS.
Time Series, Contrastive Learning, Automated Machine Learning
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3680086isbn: 979-8-4007-0436-9/24/10ccs: Computing methodologies Search methodologiesccs: Information systems Data miningccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
Time series have been collected and analyzed for plenty of real-world applications, e.g., clinical diagnosis (Eldele et al., 2021; Chen et al., 2024a), electricity forecasting (Zhou et al., 2021; Liu and Chen, 2024) and various monitoring tasks (Ren et al., 2019; Zhou et al., 2020; Jing et al., 2021b, 2024b; Wang et al., 2023). A paramount challenge for time series representation learning lies in developing an efficacious encoder capable of deriving informative embeddings, which could be readily applicable to a variety of downstream tasks, e.g., classification (Zhang et al., 2022; Nie et al., 2022; Chen et al., 2024b), forecasting (Yue et al., 2022; Jing et al., 2022c) and anomaly detection (Ren et al., 2019; Li et al., 2021; Zheng et al., 2024a).
@@
A predominant representation learning paradigm is Contrastive Learning (CL) (Liu et al., 2021; Zheng et al., 2024b), of which the core idea is to train models by pulling the positive embedding pairs closer and pushing the negative embedding pairs far apart (Jaiswal et al., 2020; Chen et al., 2020; Jing et al., 2021a; Zheng et al., 2022; Feng et al., 2022; Zheng et al., 2021). The preponderance of existing research in CL as applied to time series primarily emphasizes the development of sophisticated data augmentation techniques and the construction of contrastive pairs (Liu et al., 2023). For instance, TS-TCC (Eldele et al., 2021) employs jittering and scaling to produce varied perspectives of the input time series. These generated perspectives, originating from the same data sample, are then treated as positive instances within the model’s training regime. TS2Vec (Yue et al., 2022) uses random cropping to generate positive pairs and treat other data points in the context as negative samples. TF-C (Zhang et al., 2022) augments the input via time-domain and frequency-domain masking, and embeddings from different samples are treated as negative pairs.
@@
Despite the effectiveness of these approaches, the majority are tailored with manual construction of Contrastive Learning Strategies (CLS) specific to particular datasets and tasks, representing distinct examples within the broader spectrum of general CLS space, which could contain millions even trillions of choices. Crafting CLS for specific tasks or datasets in such a huge space through heuristic methods could be challenging, often demanding extensive domain expertise and considerable manual effort through numerous trial-and-error iterations. A natural question arises: is it possible to build a system to automatically find a suitable CLS for a given task? A pertinent question emerges: can we identify a CLS that is generally effective across a dynamic range of tasks and datasets?
@@
Aiming at automating the process of machine learning, Automated Machine Learning (AutoML) (He et al., 2021) has become popular in recent years. In the field of CL, some works optimize data augmentation strategies on images (Tian et al., 2020) and graphs (You et al., 2021; Feng et al., 2024; Yin et al., 2022). For time series, a recent work InfoTS (Luo et al., 2023) searches for optimal data augmentations based on information theory. These studies predominantly focus on data augmentation, but often overlook other crucial dimensions of CL, such as loss functions, the construction of contrastive pairs, and embedding transformations, which are typically crafted by human intervention. Specifically, a CLS encompasses diverse loss functions (e.g., InfoNCE (Oord et al., 2018) and Triplet Loss (Hoffer and Ailon, 2015)) and metric functions (e.g., dot product, cosine similarity, negative Euclidean distance) that affect representation learning quality (Zheng et al., 2022; Liu et al., 2021; Jing et al., 2022b; Yan et al., 2023; Jing et al., 2024a). Contrastive pair construction (Oord et al., 2018; Yue et al., 2022; Eldele et al., 2021; Jing et al., 2022a; Zheng et al., 2023; Li et al., 2022b) also plays a key role in distinguishing sample-level characteristics or temporal dependencies. Embedding techniques, including augmentation (Wang et al., 2019) and normalization (Wang and Isola, 2020), further influence the pretrained model performance. Few works have been conducted on jointly considering all these aspects of the contrastive learning on time series.
@@
To reduce the burden of excessive prior domain knowledge and massive manual trials of CL on time series for specific tasks, in this paper, we present an Automated Contrastive Learning (AutoCL) framework at Microsoft, which aims at learning to contrastively learn the representations for various time series tasks. Specifically, we first construct a comprehensive solution space of CLS, which not only covers the most important configurable dimensions of CL, but also considers the modest range of options for each dimension. The whole space is comprised of dimensions including data augmentations, embedding transformations, contrastive pair construction, and contrastive learning objectives, with up to $10^{12}$ level of options in total. To efficiently target the suitable CLS in such a huge solution space, we further introduce a reinforcement learning algorithm which directly optimizes the CL performance on downstream tasks in validation datasets. Compared with the recent automation-based CL method InfoTS (Luo et al., 2023), our method takes the two-fold advantage from more comprehensive solution space and the direct optimization approach on CL, while InfoTS only searches data augmentation configurations based on task-agnostic criteria, which does not consider the intrinsic connection between CLS and the downstream task, thus may result in targeting an improper CLS.
@@
We empirically evaluate AutoCL and the derived CLS over three downstream tasks including classification, forecasting, and anomaly detection on real-world time series datasets. We first directly apply AutoCL over each task to obtain its suitable CLS, which could significantly outperform existing CL methods, demonstrating the superiority of AutoCL. Then we derive a Generally Good Strategy (GGS) based on the candidate CLS found by AutoCL from several tasks and test GGS on all tasks. The results show that GGS could achieve remarkable performance for all tasks, showing its strong transferability across tasks and datasets. We also provide extensive empirical analysis on the relationship between the candidate CLS and the model’s performance on downstream tasks, hoping to provide guidance for the future design of CLS. We also tested our solution in a real-world application of epilepsy seizure detection task in Shanghai Ruijin Hospital, which has generally good performance. Our major contributions are summarized as follows:
@@
  • We introduce AutoCL to automatically derive suitable CLS for various tasks and datasets on time series. A comprehensive solution space covering the most critical dimensions of CL with an efficient reinforcement learning algorithm has been proposed.
@@ 2. Methodology
In this section, we first formulate the problem of the automated contrastive learning strategy search and then introduce a method called AutoCL for the problem. We will elaborate the two components of AutoCL: the solution space and the search algorithm.
@@ 2.1. Problem Formulation
Denote $\mathcal{D}$ as a time series dataset comprised of training $\mathcal{D}_{\text{train}}$, validation $\mathcal{D}_{\text{val}}$ and test $\mathcal{D}_{\text{test}}$ sets; $\mathcal{T}$ as a downstream task along with evaluation metrics $\mathcal{M}$; $f^{E}$ as the time series encoder. The problem of automated Contrastive Learning Strategy (CLS) search has two sub-problems: (1) define a solution space $\mathcal{A}$; (2) build a search algorithm $f$ to find a suitable CLS $A\in\mathcal{A}$ based on the validation performance on $\mathcal{D}_{\text{val}}$ regarding $\mathcal{M}$, s.t., after pre-training with $A$ on $\mathcal{D}_{\text{train}}$, $f^{E}$ will have a good performance on $\mathcal{D}_{\text{test}}$.
@@ 2.2. Solution Space
Based on the literature and our own experiences, the space is defined by following the principles (You et al., 2020): (1) encompasses the key dimensions of CLS based on prior human-designed strategies; (2) accounts for a moderate spectrum of choices for each dimension.
An illustration of a CLS is shown in Figure 1. The essence of CL is to extract semantic-preserving embeddings invariant to small data perturbations (Tian et al., 2020), and thus data augmentation is the cornerstone dimension of CLS. The commonly used practices to enhance the robustness of representation are embedding transformations, e.g., embedding augmentation (Wang et al., 2019) and normalization. There are various aspects to construct positive and negative embedding pairs, where different aspects reflect different characteristics of data. Different tasks might place emphasis on different aspects. Finally, different similarity functions and different CL losses could have divergent impacts on the learned representations. In summary, we consider four dimensions: data augmentations, embedding transformations, contrastive pair construction, and contrastive losses with around $3\times 10^{12}$ options in total. A summary of the dimensions, sub-dimensions, and their options is shown in Table 1, and we describe the details in the following content.
@@
Frequency Masking (Zhang et al., 2022)
@@
Similarity Functions Dot, Cosine, Euclidean Distance
@@
Data Augmentations. Data augmentations transform the input data into different but related views, which are the cornerstones of a CLS. We consider 6 most commonly adopted data augmentations as the sub-dimensions, including resizing (length) (Mehari and Strodthoff, 2022), rescaling (amplitude) (Han et al., 2021), jittering (Eldele et al., 2021), point masking (Zerveas et al., 2021), frequency masking (Zhang et al., 2022), and random cropping (Yue et al., 2022). Each augmentation is associated with a parameter $p\in[0,1]$. Let $\mathbf{x}\in\mathbb{R}^{T\times c}$ be an input time series, where $T$ and $c$ are the length and the number of variables. For example, $p$ in the point masking refers to the ratio of the input data points to be masked: $\mathbf{x}=\mathbf{x}\odot\mathbf{m}$, where $\mathbf{m}[t]\sim Bernoulli(1-p)$. For more details of each augmentation, please refer to the corresponding paper. We discretize the value of $p$ into 11 choices: $\{0.0,0.1,\cdots,0.8,0.9,0.95\}$. For the first value, we use $0.0$ to represent the setting where a specific data augmentation is not enabled. For the last value, we do not use $1.0$ since it is meaningless for some data augmentations in practice. For example, in point masking, $p=1.0$ means masking out all the input data.
Additionally, the order of applying data augmentations also influences the learned embeddings (Cubuk et al., 2019). For instance, the outcome of applying point masking followed by random cropping differs from first applying random cropping and then point masking. Therefore, we design 5 different orders of applying data augmentations. In summary, there are totally $5\times 11^{6}$ options for data augmentations.
Embedding Transformations. Embedding augmentations (Wang et al., 2019) and normalization (Wang and Isola, 2020) have been explored and proven useful in the literature. For embedding augmentations, we consider embedding jittering (Wang et al., 2019) and embedding masking as the sub-dimensions. Similar to the jittering and point masking in data augmentations, they are also associated with a parameter $p\in[0,1]$, whose value is also discretized into 11 values. For normalization, we consider to use no norms, $l_{2}$ norm, and layer norm (Ba et al., 2016) as the options. In summary, there are $3\times 11^{2}$ options for embedding transformations.
@@
Let $\mathbf{h}_{1},\mathbf{h}_{2}\in\mathbb{R}^{T\times d}$ be the embeddings of two views $\mathbf{x}_{1}$, $\mathbf{x}_{2}$ of an input $\mathbf{x}$, where $T$ and $d$ are the length and dimension size. The instance contrast aims to distinguish embeddings of two different time series instances $\mathbf{x},\mathbf{x}^{\prime}$ by pulling $\mathbf{h}_{1}$ and $\mathbf{h}_{2}$ (also $\mathbf{h}_{1}^{\prime}$ and $\mathbf{h}_{2}^{\prime}$) closer and pushing $\mathbf{h}_{1},\mathbf{h}_{2}$ far away from $\mathbf{h}_{1}^{\prime},\mathbf{h}_{2}^{\prime}$. The temporal contrast aims to preserve the local temporal information. For a time step $t$, it pulls $\mathbf{h}_{1}[t]$ and $\mathbf{h}_{2}[t]$ closer and optionally also $\mathbf{h}_{1}[t]$ and its temporally adjacent embeddings $\mathbf{h}_{1}[t-1],\mathbf{h}_{1}[t+1]$ closer. It pushes $\mathbf{h}_{1}[t]$ away from $\mathbf{h}_{1}[t^{\prime}]$, where $t^{\prime}\notin\{t-1,t+1\}$. These two contrasts might be sufficient for short time series, e.g., $T<100$, however, they are incapable of capturing multi-scale dynamics of long time series, e.g., $T>1000$ (Yue et al., 2022). We follow (Yue et al., 2022) and apply hierarchical pooling over $\mathbf{h}_{1},\mathbf{h}_{2}$ to obtain multi-scale embeddings $\mathbf{h}_{1}^{(s)},\mathbf{h}_{2}^{(s)}\in\mathbb{R}^{T_{s}\times d}$, where $s\in\{1,\cdots,S\}$ is the scale index and $T_{s}$ is the length. The hierarchical pooling is associated with two parameters: pooling operator (average or max pooling) and kernel size ($\{0,2,3,5\}$). The above instance contrast and temporal contrast are applied for all the scales. In addition, cross-scale contrast (Bachman et al., 2019) could further improve the cross-scale consistency. Specifically, if $\mathbf{h}^{(s+1)}[t_{s+1}]$ is pooled from $\mathbf{h}^{(s)}[t_{s}]$, then they should be pulled closer, otherwise, should be pushed away from each other.
@@
In practice, one must choose at least one contrastive aspect for CL. Based on prior studies and our experiences, the instance contrast is the most indispensable aspect, which is always included. As a result, there are 2=3-1 contrast aspects, 2 pooling operators, 4 kernel sizes and 2 options (include temporal adjacent embeddings or not) for the temporal contrast, which means we have 32 options in total.
@@
Contrastive Losses. Different types of contrastive losses behave differently. For example, InfoNCE (Oord et al., 2018) guides the model’s learning by maximizing the probabilities of the positive pairs, and Triplet loss (Hoffer and Ailon, 2015) maximizes the similarity score of the positive pairs. A contrastive loss is usually comprised of a similarity function, measuring the similarity of embeddings, and a temperature parameter (Oord et al., 2018), controlling the magnitude of similarity scores. For the similarity function, we consider dot product, cosine, and negative Euclidean distance. For the temperature parameter, we consider $\{10^{-2},10^{-1},10^{0},10^{1},10^{2}\}$. In summary, there are 2 loss types, 3 similarity functions, and 5 temperatures, resulting in 30 options.
2.3. Search Algorithm
As shown in Section 2.2, there are around $3\times 10^{12}$ total strategies in the solution space. Therefore, it could be extremely challenging and resource-consuming to manually pick a suitable CLS from the entire space. In this paper, we aim to automate the process of searching the suitable CLS within the solution space. In the following content, we first present an overview of the algorithmic part of AutoCL, and then elaborate its details, including the controller network and the two phases: candidate search and candidate evaluation.